OpenStudy (anonymous):

one pump can drain a pool in 10 min . when second pump is also used, the pool only takes 9 min to drain. How long would it take the second pump to drain if it were the only pump in use. a. 90minutes b. 89minutes c. 4 14/9 minutes d. 1/90 minutes

OpenStudy (anonymous):

a

OpenStudy (amistre64):

a is right....

OpenStudy (anonymous):

It is a. Assume we have a 10ft pool, then the first pump drains at a rate of 1ft/min in order to empty the pool in 10 minutes. Denote the first pump as A. So A = 1ft/min Now after adding a second pump say B, then A + B drains the pool in 9 minutes. So A+ B = 10ft/9mins. Now plug in A, 1ft/min + B = 10ft/9min. B = 10ft/9min - 1ft/min. B = 10ft/9min - 9ft/9min = 1ft/ 9min. Multiplying the ratio by 10 you would get 10ft/ 90mins. So It would take 90 minutes to drain 10 ft with B.

OpenStudy (amistre64):

1/10 + b = 1/9 per minute. b = 1/90 per minute. 90 minutes to drain it complete...